OpenStudy (anonymous):

Solving Equations: 5x/4 + 1/2= x - 1/2

OpenStudy (shane_b):

Do you know where to start?

OpenStudy (anonymous):

I know I have to get rid of the 1/2 on the left, right?

OpenStudy (shane_b):

If you want to start there, that's fine. What does that leave you with?

OpenStudy (anonymous):

5x/4=2(x- 1/2)?

OpenStudy (shane_b):

Close but no :) Subtract 1/2 from both sides to get:\[\frac{5x}{4}=x-\frac{1}{2}-\frac{1}{2}=x-1\]

OpenStudy (shane_b):

Do you see why the above is correct?

OpenStudy (anonymous):

yes, your subtracting 1/2 from both sides, not flipping the fraction and multiplying both sides like I did

OpenStudy (shane_b):

Ok...so now you have:\[\frac{5x}{4}=x-1\]Now what?

OpenStudy (shane_b):

I think you have it from here...let's see :)

OpenStudy (anonymous):

multiply both sides by 4 leaving me with 5x=-4 then divide by five? x=-4/5

OpenStudy (shane_b):

Check your answer...there's a mistake there

OpenStudy (shane_b):

(in the first line)

OpenStudy (anonymous):

lol divide by five... my bad

OpenStudy (shane_b):

I think you took the wrong off ramp :) \[\frac{5x}{4}=x-1\]\[5x=4(x-1)\]\[5x=4x-4\]\[x=-4\]

OpenStudy (anonymous):

ooooh, thank you so much I've been struggling with this problem since this morning

OpenStudy (shane_b):

No worries...it takes practice

OpenStudy (lgbasallote):

here comes another method... \[\frac{5x}{4} + \frac 12 = x - \frac 12\] multiply everything by the LCM..in this case 4 (so the denominator thingies will get cancelled) \[5x + 2 = 4x - 2\] subtract 4x from both sides \[x + 2 = -2\] subtract 2 from both sides \[x = -4\] although...this method takes more practice...